[04:20, 07/09/2021] Ibrahim Mohamed: Batman, whose mass is 80.0kg, is dangling on the

free end of a 12.0-m rope, the other end of which is fixed

to a tree limb above. He is able to get the rope in motion

[04:21, 07/09/2021] Ibrahim Mohamed: as only Batman knows how, eventually getting it to swing

enough that he can reach a ledge when the rope makes a

60.0°angle with the vertical. How much work was done by

the gravitational force on Batman in this maneuver?


1
Expert's answer
2021-09-07T14:07:38-0400

"g = 9.8\frac{m}{s^2}"

"l = 12m"

"m = 80kg"

"\alpha=60\degree"

"W_g= mg\Delta h"

"\Delta h = l*\cos \alpha-l= 12*\cos60\degree-12=-6m"

"W_g= 80*9.8*(-6)= -4707J"


"\text{Answer: The work of gravitational force equal }-4707\ J."